# On the Operational Resilience of CBDC: Threats and Prospects of Formal Validation for Offline Payments

### v1

@@ 3 A Methodology Based on Formal Methods
As far as economics and finance are concerned, the application of formal methods is rather limited. In the case of traditional centralized systems, we are aware of very few case studies; among them we mention [4, 53, 69, 64, 55, 70, 72]. In the case of modern decentralized systems, apart from some works about blockchain consensus like [78, 79, 16], the formal methods literature is focusing on smart contracts [76]; see, e.g., [13].
@@ 3.3 An Illustrative Example
Operator $P_{1}\mathop{\|_{S}}P_{2}$ describes the fact that processes $P_{1}$ and $P_{2}$ run in parallel and have to synchronize when executing actions in $S$, which is empty in the first equation above. Note that $\hbox{}\mathord{\it Prod}$ (resp. $\hbox{}\mathord{\it Cons}$) repeatedly tries to deposit items into (resp. withdraw items from) the buffer.
@@
It turns out that $\hbox{}\mathord{\it PC}_{{\rm conc},2}$ is strongly bisimilar [58] to $\hbox{}\mathord{\it ProdCons}_{0/2}$, i.e., they are able to mimic each other’s behavior stepwise. The bisimulation proving this fact has been represented graphically by giving the same color to states in the same equivalence class and different colors to different equivalence classes. The depicted relation is a strong bisimulation because in both automata:
@@
It turns out that $\hbox{}\mathord{\it PC}_{{\rm pipe},2}$ is weakly bisimilar [58] to $\hbox{}\mathord{\it ProdCons}_{0/2}$, in the sense that $\tau$-actions are ignored in the bisimulation game. The weak bisimulation proving this fact has been represented graphically by means of colors. The depicted relation is a weak bisimulation because in both automata:
@@ 4 Conclusions
On the other hand, the impossibility results described in Section 1.3, concerning the inability in general to theoretically guarantee the quality and correctness of the protocols and the software required to implement them, mean that we cannot absolutely ensure the perfect and secure functioning of the CBDC infrastructure. This leaves the risk of possible trapdoors, open to exploitation by bad actors, which could be used to create counterfeit transactions and money.

### v2

@@ 3 A Methodology Based on Formal Methods
As far as economics and finance are concerned, the application of formal methods is rather limited. In the case of traditional centralized systems, we are aware of very few case studies; among them we mention [And_1999, LRK_2013, SIR_2013, PI_2017, MMNOSV_2018, SS_2019, Sto_2023]. In the case of modern decentralized systems, apart from some works about blockchain consensus like [VYC_2022, Ves_2023, BEFR_2025], the formal methods literature is focusing on smart contracts [Sza94]; see, e.g., [BZ_2018].
@@ 3.3 An Illustrative Example
Operator $P_{1}\mathop{\|_{S}}P_{2}$ describes the fact that processes $P_{1}$ and $P_{2}$ run in parallel and have to synchronize when executing actions in $S$; this set is empty in the first equation above in the case of the two positions of the buffer. Note that $\hbox{}\mathord{\it Prod}$ (resp. $\hbox{}\mathord{\it Cons}$) repeatedly tries to deposit items into (resp. withdraw items from) the buffer.
@@
It turns out that $\hbox{}\mathord{\it PC}_{{\rm conc},2}$ is strongly bisimilar [Mil89a] to $\hbox{}\mathord{\it ProdCons}_{0/2}$, i.e., they are able to mimic each other’s behavior stepwise. The bisimulation proving this fact has been represented graphically by giving the same color to states in the same equivalence class and different colors to different equivalence classes. The depicted relation is a strong bisimulation because in both automata:
@@
It turns out that $\hbox{}\mathord{\it PC}_{{\rm pipe},2}$ is weakly bisimilar [Mil89a] to $\hbox{}\mathord{\it ProdCons}_{0/2}$, in the sense that $\tau$-actions are ignored in the bisimulation game. The weak bisimulation proving this fact has been represented graphically by means of colors. The depicted relation is a weak bisimulation because in both automata:
@@ 4 Conclusions
On the other hand, the impossibility results mentioned in Section 1.3, concerning the inability in general to theoretically guarantee the quality and correctness of the protocols and the software required to implement them, mean that we cannot absolutely ensure the perfect and secure functioning of the CBDC infrastructure. This leaves the risk of possible trapdoors, open to exploitation by bad actors, which could be used to create counterfeit transactions and money.